# Does Your Multi-modal LLM Truly See the Diagrams in Visual Math Problems?

### v1

@@ 3.2 Experimental Analysis
Figure 7: A Typical Visual Perception Error by our CoT Evaluation Strategy. The example is an output from Gemini-Pro team2023gemini , where the correct reasoning of the Text-only version is distracted by the visual perception error within the diagram.
@@
In addition to foundational visual instruction-following datasets, both SPHINX-MoE and InternLM-XComposer2 extend their training regimes to include specialized mathematical problems that are either text-only or visual, such as MathQA amini2019mathqa , Geometry3K Lu2021InterGPSIG , and MathInstruct yue2023mammoth . This approach of math-specific tuning contributes to their leading performance in MathVerse. Furthermore, G-LLaVA fine-tunes LLaVA-1.5 by a large-scale visual geometric dataset containing 170K enriched problems. This targeted refinement can improve several fields (‘Len’, ‘Angle’, and ‘Apply’) within the plane geometry subject. However, since G-LLaVA’s fine-tuning data does not include problems of analytic geometry, solid geometry, and functions, it harms the related results of LLaVA-1.5 due to catastrophic forgetting, e.g., -3.5% in ‘Anal’, -5.6% in ‘Solid Geometry’, and -5.6% in ‘Functions’. This phenomenon underscores the critical role of developing extensive, high-quality visual math data for effectively training MLLMs.
@@ 3.3 Error Analysis
Figure 9: Distribution of GPT-4V’s openai2023gpt4v Errors in Reasoning and Answers. For the six problem versions in MathVerse, we provide the statistics of errors made by GPT-4V based on their occurrence in answers (‘Ans.’) and reasoning processes (‘Rea.’).
To delve into the fine-grained predictions, we select the best-performing MLLM, GPT-4V openai2023gpt4v , to understand its modes of success and failure. Our proposed CoT evaluation strategy has produced a detailed assessment of model output, including step-wise scores and explanation, reducing extensive manual effort in identifying and analyzing errors. We conduct our analysis on the two-step output from the CoT evaluation across the entire dataset, focusing on two key dimensions.
@@
Figure 10: Distribution of GPT-4V’s openai2023gpt4v Errors within Different Types. We present the statistics of four error types by GPT-4V in the six problem versions, i.e., Visual Perception Error, Reasoning Error, Calculation Error, and Knowledge Error.
@@ Appendix A Related Work
building upon the prevalence of Large Language Models (LLMs) touvron2023llama ; touvron2023llama2 ; OpenAI2023ChatGPT ; albert24mixtral ; brown2020language and large vision models Radford2021LearningTV ; kirillov2023segment ; zhang2023personalize ; zhang2023prompt ; zhang2023learning , have become increasingly prominent in the field. They extend LLMs to tackle a diverse range of tasks and domains, including the mainstream 2D images li2022blip ; instructblip ; alayrac2022flamingo ; li2023mimic and other modalities, such as 3D point clouds guo2023point ; xu2023pointllm ; hong20243d , audio han2023imagebind ; su2023pandagpt , and video zhang2023video ; chen2023videollm . Noteworthy examples like OpenAI’s GPT-4V openai2023gpt4v and Google’s Gemini team2023gemini exhibit exceptional visual understanding and reasoning capabilities, setting new benchmarks in multi-modal performance. However, their closed-source nature poses a barrier to the broader application and development of MLLMs. Concurrently, another line of work is dedicated to exploring advanced MLLMs open-source to the community. Prior efforts like LLaMA-Adapter zhang2024llamaadapter ; gao2023llamaadapterv2 , LLaVA liu2023llava ; liu2024llavanext ; liu2023improvedllava , and MiniGPT-4 zhu2023minigpt ; chen2023minigpt leverage a frozen CLIP Radford2021LearningTV model for image encoding, and inject the visual cues into LLaMA touvron2023llama for multi-modal instruction tuning. The subsequent mPLUG-Owl ye2023mplugowl ; ye2023mplugowl2 , Qwen-VL bai2023qwen , InternLM-XComposer dong2024internlm , and SPHINX lin2023sphinx ; gao2024sphinx further push the frontier of MLLMs in understanding and generalizing across visual contexts. Despite comprehensive benchmarks fu2023mme ; liu2023mmbench ; Li2023SEEDBenchBM ; xu2023lvlm on general visual instruction-following scenarios, the specific potential of MLLMs for visual mathematical problem-solving remains under-explored. In this paper, we introduce the MathVerse benchmark to comprehensively evaluate the visual mathematical reasoning and diagram understanding skills of MLLMs, providing unique perspectives for future research directions.
@@
have emerged as a significant area of focus, posing considerable challenges for large foundational models, e.g., LLMs and MLLMs. Initially, datasets in this realm are designed to address basic algebraic hendrycksmath2021 and arithmetic Roy2016SolvingGA word problems, which are relatively limited in scope and volume. Subsequent efforts, including MATH hendrycksmath2021 , GSM8K cobbe2021training , and MMLU hendryckstest2021 , expand the range and quality of textual mathematical problems. These datasets feature a broader spectrum of difficulties, establishing a robust benchmark for the evaluation of general and math-specific LLMs zhou2023solving ; yue2023mammoth ; wang2024mathcoder ; gao2023g ; luo2023wizardmath . Besides the text-only assessment, there is a growing demand for comparable, high-quality benchmarks for evaluating mathematical problem-solving in visual contexts, with the rapid progress of MLLMs. There are prior attempts, such as GeoQA Chen2021GeoQAAG , UniGeo Chen2022UniGeoUG , and Geometry3K Lu2021InterGPSIG , which focused exclusively on geometric problems. The recently proposed MathVista Lu2023MathVistaEM broadens the scope to incorporate a variety of multi-modal tasks involving mathematical reasoning, and MMMU yue2023mmmu covers college-level questions demanding intricate, domain-specific knowledge. However, our analysis identifies three main shortcomings within the current visual math benchmarks, as elaborated in Section 1 of the main paper. Therefore, we propose MathVerse specialized in the multi-modal mathematical evaluation of MLLMs, comprising twelve subjects, six problem versions, and 20K test samples. Our objective is to thoroughly investigate whether and how much MLLMs genuinely interpret visual diagrams for mathematical reasoning.
@@
Model Release Time
Source
ChatGPT ouyang2022training 2022-11
https://platform.openai.com/
GPT-4 OpenAI2023GPT4TR 2023-03
https://platform.openai.com/
Qwen-VL-Plus bai2023qwen 2023-11
https://help.aliyun.com/zh/dashscope/developer-reference/vl-plus-quick-start
Gemini-Pro team2023gemini 2023-12
https://ai.google.dev/
Qwen-VL-Max bai2023qwen 2024-01
https://help.aliyun.com/zh/dashscope/developer-reference/vl-plus-quick-start
GPT-4V openai2023gpt4v 2023-09
https://platform.openai.com/
LLaMA-Adapter V2 gao2023llamaadapterv2 2023-04
https://github.com/OpenGVLab/LLaMA-Adapter/tree/main/llama_adapter_v2_multimodal7b
LLaVA-1.5 liu2023improvedllava 2023-10
https://huggingface.co/liuhaotian/llava-v1.5-13b
MiniGPT-v2 chen2023minigpt 2023-10
https://github.com/Vision-CAIR/MiniGPT-4
mPLUG-Owl2 ye2023mplugowl2 2023-11
https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b
G-LLaVA gao2023g 2023-12
https://github.com/pipilurj/G-LLaVA/tree/main
ImageBind-LLM han2023imagebind 2023-05
https://github.com/OpenGVLab/LLaMA-Adapter/tree/main/imagebind_LLM
ShareGPT4V Chen2023ShareGPT4VIL 2023-11
https://huggingface.co/Lin-Chen/ShareGPT4V-13B
SPHINX-Plus lin2023sphinx 2023-11
https://huggingface.co/Alpha-VLLM/LLaMA2-Accessory/tree/main/finetune/mm/SPHINX/SPHINX-v2-1k
LLaVA-NeXT liu2024llavanext 2024-01
https://huggingface.co/liuhaotian/llava-v1.6-vicuna-13b
SPHINX-MoE gao2024sphinx 2024-01
https://huggingface.co/Alpha-VLLM/LLaMA2-Accessory/tree/main/finetune/mm/SPHINX/SPHINX-MoE
InternLM-XComposer2 dong2024internlm 2024-01
https://huggingface.co/internlm/internlm-xcomposer2-vl-7b
@@ Appendix B Additional Experimental Details
Free-form Question
Please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
@@
Multiple-choice Question
Please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
@@
Free-form Question
According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
Multiple-choice Question
According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
@@
Figure 11: Manual Annotations for Function Problems in MathVerse. We provide detailed annotations, e.g., function expression and properties, for the diagrams of 534 function problems, which benefits the accuracy of GPT-4V openai2023gpt4v for CoT evaluation.
@@
Our proposed CoT evaluation contains two steps, i.e., key-step extraction and multi-step scoring, which prompt GPT-4 OpenAI2023GPT4TR and GPT-4V openai2023gpt4v , respectively. The input configuration is listed in Table 7. We utilize the text-only GPT-4 in the first step to extract multiple key steps within the model’s unstructured output, without feeding the question information. In the second step, we input the extracted key-step reasoning and all the available content related to the problem into GPT-4V, allowing for a holistic assessment, including diagram interpretation, logical reasoning, and numerical computation. In Figure 11, we showcase the manual annotation for critical information within functional diagrams, e.g., function expression and properties. This assists GPT-4V in evaluating the visual perception accuracy of MLLMs for function graphs.
Table 7: Configuration for the CoT Evaluation Strategy. We conduct two evaluation phases, respectively by prompting the text-only GPT-4 OpenAI2023GPT4TR and GPT-4V openai2023gpt4v . The symbol ‘XXX’ denotes the given one-shot sample, abbreviated for brevity. The ‘Annotation’ (represented in grey) in the second phase is only required for function problems.
@@
Key-step Extraction (GPT-4) Model Output
I will give you a detailed solving procedure or a single answer for a math problem.
If it is a procedure, you need to extract the key solution steps and list them accordingly in markdown syntax. If it is just a single answer, output the answer directly.
Here are examples:
– Model output: XXX
– Extracted: 1. XXX 2. XXX 3. XXX
– Model output: 2.2
– Extracted: The single answer is 2.2
Here is what you need to extract:
Model output: $\{model\ output\}$
Extracted:
Multi-step Scoring (GPT-4V) Extracted Steps Question Diagram Answer Annotation
I will first give you a visual math problem, including the question, diagram, ground-truth answer, and detailed annotation of the diagram, and then give you a model output containing multiple key solution steps.
Please think step by step and output the Average score, along with the Final answer score in the end, as described below:
– Average score: Evaluate, based on the given question, answer, diagram, and diagram annotation, whether each solution step is correct in logical reasoning, visual perception, and numerical computation, with an incorrect score of 0 and a correct score of 1. Then, calculate the average score of multiple steps.
– Final answer score: Match the model’s final answer with the ground truth answer, scoring 1 if it matches and 0 if it doesn’t.
– If the model output only includes a single step or answer, the Average score and Final answer score are the same.
Question: $\{question\}$
Ground-truth answer: $\{answer\}$
Diagram annotation: $\{annotation\}$
Model output: $\{extracted\ steps\}$
Average score:
Final answer score:
@@
[…]
@@ C.1 Data Curation
[…]
@@ C.3 Detailed Statistics of MathVerse
Table 9: Length of Different Problem Versions in MathVerse.
@@ Appendix D Comparison to Current Benchmarks
In this section, we offer a detailed comparison between MathVerse and existing multi-modal mathematical benchmarks, i.e., geometry-specific benchmarks chen2021geoqa ; cao2022augmented ; chen2022unigeo ; lu2021inter ; seo2015solving , MathVista Lu2023MathVistaEM , and MMMU yue2023mmmu , from the following four aspects:
@@
Figure 18: Diagram Examples of Math-related Tasks in MathVista Lu2023MathVistaEM . These tasks are not strongly correlated to the mathematical reasoning skills of MLLMs, probably skewing the assessment emphasis towards visual math problems.
Figure 19: Diagram Examples with Required Knowledge in MMMU yue2023mmmu . These math problems demand MLLMs to comprehend college-level domain knowledge, potentially hindering them from fully exerting mathematical reasoning skills.
@@
Benchmark
GEOS
Geo3K
GeoQA+
MathVista
MMMU-Math
MathVerse
Test Samples
119
601
755
6,141
540
15,672
@@ Appendix E Limitation and Future Work
We have categorized the math problems in MathVerse by various criteria, including subjects, subfields, and versions featuring differing degrees of multi-modal content. These categorization approaches evaluate the capabilities of MLLMs from multiple dimensions. Nevertheless, it is also meaningful to further divide the problems based on their difficulty levels, akin to MATH hendrycksmath2021 , a text-only benchmark defining five levels of difficulty. This additional layer of differentiation can provide deeper insights into the problem-solving abilities of MLLMs across a spectrum of challenges, which we leave as future work.
@@
Moreover, the problems in MathVerse and other current mathematical benchmarks are mainly in English. Given that some multilingual MLLMs han2023imagebind ; bai2023qwen have been developed, existing evaluation cannot reveal their full capabilities when confined to a single language. The incorporation of multilingual visual math problems would not only extend the dataset’s global applicability, but also enhance the assessment of MLLMs for linguistic diversity and understanding.
@@ Appendix F Qualitative Examples
To ease the understanding, we offer a variety of qualitative examples in MathVerse. In Section F.1, we showcase the meticulously transformed six versions of visual math problems. In Section F.2, we compare the response of different MLLMs on Text-lite problems, including GPT-4V openai2023gpt4v , LLaVA-NeXT liu2024llavanext , and SPHINX-MoE gao2024sphinx . Specifically, we present the key-step extraction output by the CoT evaluation, and mark the multi-step scoring results aside. In Section F.3, we provide the response comparison of GPT-4V for three problem versions in MathVerse, i.e., Text Dominant, Text Lite, and Vision Dominant.

### v2

@@ 3.2 Experimental Analysis
Figure 7: A Typical Visual Perception Error by our CoT Evaluation Strategy. The example is an output from Gemini-Pro [22], where the correct reasoning of the Text-only version is distracted by the visual perception error within the diagram.
@@
In addition to foundational visual instruction-following datasets, both SPHINX-MoE and InternLM-XComposer2 extend their training regimes to include specialized mathematical problems that are either text-only or visual, such as MathQA [2], Geometry3K [43], and MathInstruct [64]. This approach of math-specific tuning contributes to their leading performance in MathVerse. Furthermore, G-LLaVA fine-tunes LLaVA-1.5 by a large-scale visual geometric dataset containing 170K enriched problems. This targeted refinement can improve several fields (‘Len’, ‘Angle’, and ‘Apply’) within the plane geometry subject. However, since G-LLaVA’s fine-tuning data does not include problems of analytic geometry, solid geometry, and functions, it harms the related results of LLaVA-1.5 due to catastrophic forgetting, e.g., -3.5% in ‘Anal’, -5.6% in ‘Solid Geometry’, and -5.6% in ‘Functions’. This phenomenon underscores the critical role of developing extensive, high-quality visual math data for effectively training MLLMs.
@@ 3.3 Error Analysis
Figure 9: Distribution of GPT-4V’s [47] Errors in Reasoning and Answers. For the six problem versions in MathVerse, we provide the statistics of errors made by GPT-4V based on their occurrence in answers (‘Ans.’) and reasoning processes (‘Rea.’).
To delve into the fine-grained predictions, we select the best-performing MLLM, GPT-4V [47], to understand its modes of success and failure. Our proposed CoT evaluation strategy has produced a detailed assessment of model output, including step-wise scores and explanation, reducing extensive manual effort in identifying and analyzing errors. We conduct our analysis on the two-step output from the CoT evaluation across the entire dataset, focusing on two key dimensions.
@@
Figure 10: Distribution of GPT-4V’s [47] Errors within Different Types. We present the statistics of four error types by GPT-4V in the six problem versions, i.e., Visual Perception Error, Reasoning Error, Calculation Error, and Knowledge Error.
@@ Appendix A Related Work
building upon the prevalence of Large Language Models (LLMs) [55, 56, 45, 28, 4] and large vision models [49, 29, 68, 67, 69], have become increasingly prominent in the field. They extend LLMs to tackle a diverse range of tasks and domains, including the mainstream 2D images [35, 15, 1, 31] and other modalities, such as 3D point clouds [23, 60, 27], audio [24, 52], and video [65, 6]. Noteworthy examples like OpenAI’s GPT-4V [47] and Google’s Gemini [22] exhibit exceptional visual understanding and reasoning capabilities, setting new benchmarks in multi-modal performance. However, their closed-source nature poses a barrier to the broader application and development of MLLMs. Concurrently, another line of work is dedicated to exploring advanced MLLMs open-source to the community. Prior efforts like LLaMA-Adapter [66, 20], LLaVA [39, 38, 37], and MiniGPT-4 [72, 11] leverage a frozen CLIP [49] model for image encoding, and inject the visual cues into LLaMA [55] for multi-modal instruction tuning. The subsequent mPLUG-Owl [61, 62], Qwen-VL [3], InternLM-XComposer [16], and SPHINX [36, 21] further push the frontier of MLLMs in understanding and generalizing across visual contexts. Despite comprehensive benchmarks [17, 40, 33, 59] on general visual instruction-following scenarios, the specific potential of MLLMs for visual mathematical problem-solving remains under-explored. In this paper, we introduce the MathVerse benchmark to comprehensively evaluate the visual mathematical reasoning and diagram understanding skills of MLLMs, providing unique perspectives for future research directions.
@@
have emerged as a significant area of focus, posing considerable challenges for large foundational models, e.g., LLMs and MLLMs. Initially, datasets in this realm are designed to address basic algebraic [26] and arithmetic [50] word problems, which are relatively limited in scope and volume. Subsequent efforts, including MATH [26], GSM8K [14], and MMLU [25], expand the range and quality of textual mathematical problems. These datasets feature a broader spectrum of difficulties, establishing a robust benchmark for the evaluation of general and math-specific LLMs [71, 64, 57, 19, 44]. Besides the text-only assessment, there is a growing demand for comparable, high-quality benchmarks for evaluating mathematical problem-solving in visual contexts, with the rapid progress of MLLMs. There are prior attempts, such as GeoQA [10], UniGeo [8], and Geometry3K [43], which focused exclusively on geometric problems. The recently proposed MathVista [41] broadens the scope to incorporate a variety of multi-modal tasks involving mathematical reasoning, and MMMU [63] covers college-level questions demanding intricate, domain-specific knowledge. However, our analysis identifies three main shortcomings within the current visual math benchmarks, as elaborated in Section 1 of the main paper. Therefore, we propose MathVerse specialized in the multi-modal mathematical evaluation of MLLMs, comprising twelve subjects, six problem versions, and 20K test samples. Our objective is to thoroughly investigate whether and how much MLLMs genuinely interpret visual diagrams for mathematical reasoning.
@@ Appendix B Additional Experimental Details
Free-form Question Please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
@@
Multiple-choice Question Please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
@@
Free-form Question According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the final value, e.g., 1, 2.5, 300, at the end.
Multiple-choice Question According to the question shown in the image, please first conduct reasoning, and then answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.
@@
Figure 11: Manual Annotations for Function Problems in MathVerse. We provide detailed annotations, e.g., function expression and properties, for the diagrams of 534 function problems, which benefits the accuracy of GPT-4V [47] for CoT evaluation.
@@
Our proposed CoT evaluation contains two steps, i.e., key-step extraction and multi-step scoring, which prompt GPT-4 [46] and GPT-4V [47], respectively. The input configuration is listed in Table 7. We utilize the text-only GPT-4 in the first step to extract multiple key steps within the model’s unstructured output, without feeding the question information. In the second step, we input the extracted key-step reasoning and all the available content related to the problem into GPT-4V, allowing for a holistic assessment, including diagram interpretation, logical reasoning, and numerical computation. In Figure 11, we showcase the manual annotation for critical information within functional diagrams, e.g., function expression and properties. This assists GPT-4V in evaluating the visual perception accuracy of MLLMs for function graphs.
Table 7: Configuration for the CoT Evaluation Strategy. We conduct two evaluation phases, respectively by prompting the text-only GPT-4 [46] and GPT-4V [47]. The symbol ‘XXX’ denotes the given one-shot sample, abbreviated for brevity. The ‘Annotation’ (represented in grey) in the second phase is only required for function problems.
@@
Key-step Extraction (GPT-4) Model Output I will give you a detailed solving procedure or a single answer for a math problem. If it is a procedure, you need to extract the key solution steps and list them accordingly in markdown syntax. If it is just a single answer, output the answer directly. Here are examples: – Model output: XXX – Extracted: 1. XXX 2. XXX 3. XXX – Model output: 2.2 – Extracted: The single answer is 2.2 Here is what you need to extract: Model output: $\{model\ output\}$ Extracted:
Multi-step Scoring (GPT-4V) Extracted Steps Question Diagram Answer Annotation I will first give you a visual math problem, including the question, diagram, ground-truth answer, and detailed annotation of the diagram, and then give you a model output containing multiple key solution steps. Please think step by step and output the Average score, along with the Final answer score in the end, as described below: – Average score: Evaluate, based on the given question, answer, diagram, and diagram annotation, whether each solution step is correct in logical reasoning, visual perception, and numerical computation, with an incorrect score of 0 and a correct score of 1. Then, calculate the average score of multiple steps. – Final answer score: Match the model’s final answer with the ground truth answer, scoring 1 if it matches and 0 if it doesn’t. – If the model output only includes a single step or answer, the Average score and Final answer score are the same. Question: $\{question\}$ Ground-truth answer: $\{answer\}$ Diagram annotation: $\{annotation\}$ Model output: $\{extracted\ steps\}$ Average score: Final answer score:
@@
1https://homework.study.com2https://www.ixl.com/math3https://mathspace.co/us4https://mathpix.com
@@ C.1 Data Curation
We comprehensively collect visual math problems from existing datasets [43, 9, 51] and public question repositories1,2,3. We specifically select high-quality plane geometric problems from current benchmarks, which showcase various question types, moderate question length, diverse diagram styles, and appropriate solving difficulty. For the manually collected problems of three subjects (plane geometry, solid geometry, and functions), we apply the Mathpix tool4 to accurately extract the question texts, diagrams, explanations, and answers from the website. We strictly comply with copyright and licensing rules, ensuring that we refrain from using data from sites that forbid copying and redistribution. After the initial collection, we obtain around 3.5K visual math problems, with 1.5K from existing datasets and 2K newly collected.
@@ Appendix D Comparison to Current Benchmarks
In this section, we offer a detailed comparison between MathVerse and existing multi-modal mathematical benchmarks, i.e., geometry-specific benchmarks [9, 5, 7, 42, 51], MathVista [41], and MMMU [63], from the following four aspects:
@@
Figure 18: Diagram Examples of Math-related Tasks in MathVista [41]. These tasks are not strongly correlated to the mathematical reasoning skills of MLLMs, probably skewing the assessment emphasis towards visual math problems.
Figure 19: Diagram Examples with Required Knowledge in MMMU [63]. These math problems demand MLLMs to comprehend college-level domain knowledge, potentially hindering them from fully exerting mathematical reasoning skills.
@@
Benchmark GEOS Geo3K GeoQA+ MathVista MMMU-Math MathVerse
Test Samples 119 601 755 6,141 540 15,672
@@ Appendix E Limitation and Future Work
We have categorized the math problems in MathVerse by various criteria, including subjects, subfields, and versions featuring differing degrees of multi-modal content. These categorization approaches evaluate the capabilities of MLLMs from multiple dimensions. Nevertheless, it is also meaningful to further divide the problems based on their difficulty levels, akin to MATH [26], a text-only benchmark defining five levels of difficulty. This additional layer of differentiation can provide deeper insights into the problem-solving abilities of MLLMs across a spectrum of challenges, which we leave as future work.
@@
Moreover, the problems in MathVerse and other current mathematical benchmarks are mainly in English. Given that some multilingual MLLMs [24, 3] have been developed, existing evaluation cannot reveal their full capabilities when confined to a single language. The incorporation of multilingual visual math problems would not only extend the dataset’s global applicability, but also enhance the assessment of MLLMs for linguistic diversity and understanding.
@@ Appendix F Qualitative Examples
To ease the understanding, we offer a variety of qualitative examples in MathVerse. In Section F.1, we showcase the meticulously transformed six versions of visual math problems. In Section F.2, we compare the response of different MLLMs on Text-lite problems, including GPT-4V [47], LLaVA-NeXT [38], and SPHINX-MoE [21]. Specifically, we present the key-step extraction output by the CoT evaluation, and mark the multi-step scoring results aside. In Section F.3, we provide the response comparison of GPT-4V for three problem versions in MathVerse, i.e., Text Dominant, Text Lite, and Vision Dominant.